OpenStudy (anonymous):

You have a jar of 200 jelly beans (yum!), 15 are red. -What are the chances of grabbing three red jellybeans at once (one grab, not three) -What are the chances of grabbing four red jelly beans and one non-red at once (again, one grab ,not five) Thanks!

OpenStudy (anonymous):

three at once \[\frac{15}{200}\times \frac{14}{199}\times \frac{13}{198}\]

OpenStudy (anonymous):

not sure how "one grab" is any different than three. like the difference between rolling on die three times or three dice once. (no difference)